Nerve impulses travel at different speeds. Touch receptors travel at 76.2 m/s, pain receptors travel at 0.610 m/s. I'm 6' 5" tall (1" = 2.540cm) and if I stub my toe. what is the time delay between hittimg my toe and feeling it? What is the time delay between feeling it and feeling the pain?
1
Expert's answer
2013-02-25T11:54:38-0500
6' 5" = 77 inches
77/39.4 = 1.96 metres

the time delay between hittig toe and feeling it:
t_f = h/v_t
h = 1.96 metres
v_t - speed of touch receptors
t = 1.96/76.2 = 0.026 sec

the time delay between feeling it and feeling the pain:
t= t_p - t_f,
t_p = h/v_p
v_p - speed of pain receptors
t = h ( 1/v_p - 1/v_f ) = 1.96 ( 1/0.61 - 1/76.2) = 3.2 sec